# Zero-Shot Skeleton-based Action Recognition with Dual Visual-Text Alignment

### v1

@@ III-B2 Augmented Alignment
TABLE I: Comparison of state-of-the-art methods with DVTA on the NTU-60 and NTU-120 datasets.
@@
TABLE II: Optimized experimental setting on NTU-60, NTU-120, and PKU-MMD datasets.
@@
IV Experiments
IV-A Datasets
IV-B Implementation Details
For data preprocessing, we resize the skeleton sequence to 50 frames using linear interpolation and removal of invalid frames, which is the same as on Cross-CLR [48]. For text modality, 768-dimensional word embeddings are obtained by a pre-trained text feature extractor from Sentence-Bert [49]. For skeleton modality, Shift-GCN [25] are used as feature encoders to obtain 256-dimensional features. The dimensionality is then increased to 768 dimensions by the nonlinear feature projector. The semantic features are passed through a linear layer to enhance the learning of similarity and to retain as much semantic information as possible. The optimizer employed is Adam, with a learning rate of 1e-5, and the training process spans 100 epochs, utilizing the CosineAnnealing scheduler. The mini-batch size is 128, and the temperature parameter for the contrastive loss is set to 0.1. Our implementation relies on the PyTorch framework, and all experiments are carried out using an NVIDIA GeForce RTX 4090.
IV-C Experimental Results
TABLE III: Influence of different modules on the NTU-60 and NTU-120 datasets.
@@ IV-C Experimental Results
TABLE IV: Effect on different settings of Direct Alignment module.
@@
TABLE V: Effect on different settings of the hyperparameter $\gamma$ on Augmented Alignment module.
@@
Comparison with State-of-the-arts: We compare the proposed DVTA with the state-of-the-art approaches to demonstrate its effectiveness in zero-shot learning. To ensure a fair comparison, our experimental setup strictly remains consistent with that of SynSE [35], including the use of identical skeleton feature and semantic feature encoders as well as the same split between seen and unseen classes. We also refrain from utilizing Semantic Descriptions Enhancement module. We employ Shift-GCN [25] as the skeleton feature extractor and leverage the pre-trained Sentence-Bert [49] for semantic feature extraction. Follwing SynSE [35], we establish two distinct category divisions for each dataset. The NTU-60 dataset divided its 60 categories into 5 and 12 unseen classes, corresponding to the 55/5 and 48/12 splits in the table. For the NTU-120 dataset, which boasts a larger sample size and more categories, the number of unseen classes increased to 10 and 24, resulting in the 110/10 and 96/24 splits. Table I presents a comparison of the performance of the DVTA with state-of-the-art methods on the NTU-60 and NTU-120 datasets. The proposed DVTA achieves excellent performance in zero-shot learning. Compared with SMIE, DVTA achieves 1.3% and 3.96% performance improvement on the 55/5 split and 48/12 split of the NTU-60 dataset, respectively. For the NTU-120 dataset, DVTA achieves 9.15% and 6.51% on the 110/10 split and 96/24 split, respectively. “DVTA w/o SDE” means zero-shot learning on the pure network architecture without Semantic Description Enhancement. The performance of all settings exceeds the SMIE, reflecting the superiority of our network architecture. It is noteworthy that as the number of unseen classes increases, zero-shot learning becomes more challenging. However, our DVTA still performs remarkably well, especially on the more challenging 110/10 split and 96/24 split, which fully demonstrates the significant advantage of our model in handling scenarios with more unseen classes.
Optimized Experimental Setting Results: To better validate the effectiveness of our method in zero-shot learning, we conduct comparison experiments under an optimized setting [9]. We use three large-scale skeleton datasets, i.e., NTU-60, NTU-120, and PKU-MMD as benchmark datasets. For each dataset, we conduct three randomized splits and average the results to reduce biases introduced by different class splits. The seen/unseen class splits are set as 55/5, 110/10 and 46/5, respectively, following the class split settings in SMIE [9]. We use the classic ST-GCN [22] as the visual feature extractor and Sentence-BERT [49] to extract semantic feature. The comparison results are shown in Table II. Our method outperforms others across all three datasets, achieving improvements of 10.46%, 3.96%, and 9.91%, respectively. The results indicate that DVTA effectively aligns the skeleton and text modalities, and the dual-alignment approach is more conducive to zero-shot learning.
IV-D Ablation Study and Visualizations
Influence of Modules: To evaluate the effectiveness of each module in the Dual Visual-Text Alignment (DVTA), we conduct a systematic ablation study. The experiments are performed under the optimized settings, with three split tests on the NTU-60 and NTU-120 datasets, and the average performance is reported, as shown in Table III.
The DA module can be applied as an independent model for zero-shot skeleton action recognition, optimized through training with the $L_{1}$ loss function. This module comprises a deep nonlinear mapping projector for skeleton features and a linear layer for text features, allowing for the investigation of the utility of these components through ablation experiments. The results of the ablated experiments on the DA module’s components are shown in Table IV, where “Semantic” denotes the use of the text mapping, and “Visual” denotes the use of the designed deep nonlinear mapping projector for skeleton features. The results indicate that using the complete DA module more effectively reduces the gap between modalities, thereby improving alignment quality.
TABLE VI: Ablation on Semantic Descriptions Enhancement module.
@@ IV-D Ablation Study and Visualizations
TABLE VII: Experiments with the loss function on the NTU-60 dataset with different temperature parameters.
@@
Effect of Augmented Alignment module: In the Augmented Alignment (AA) module, we design a LeakySigmoid function to convert the output value of the network $E$ into a similarity score $G$. It has a hyperparameter, $\gamma$, which controls the slope of the function at negative values. To determine the optimal choice of $\gamma$ and demonstrate the effectiveness of the LeakySigmoid function, we conducted experiments on all four splits based on the complete model. The experimental results are shown in Table V, where “None” indicates no utilization of any transformation function, and “Sigmoid” signifies the application of the Sigmoid function to output similarity scores. Through comparison, it is found that the performance of the LeakySigmoid function consistently surpasses that of the Sigmoid function, regardless of the value set for the $\gamma$ hyperparameter. This discovery effectively validates the usefulness of the LeakySigmoid function in the AA module. Specifically, when $\gamma$ was set to 0.01, the model exhibits the best performance on two splits. To ensure consistency and comparability in our experiments, we uniformly set the $\gamma$ hyperparameter to 0.01 throughout the entire experimental process.
Effect of Semantic Descriptions Enhancement module: We encode the action text using different text encoders. Table 5 presents the SDE ablation study results using Sentence-BERT [49] and CLIP [50] as semantic feature extractors on the NTU-60 and NTU-120 datasets. The results indicate that the use of the Semantics Description Enhancement (SDE) module leads to significant performance improvements across all split tests, regardless of the encoder used. For the NTU-120 dataset, under the 96/24 split, the accuracy of action labels and the accuracy when enhanced with action description encoding are 48.53% and 51.81%, respectively, using the BERT encoder. Figure 6 shows the per-class accuracy changes after applying the SDE module. The visualization reveals substantial improvements in the accuracy of certain hand-related action categories, such as “waving”, “picking up”, and “punching and kicking”.
Effect of the Loss Functions: We evaluate the InfoNCE, SoftmaxCE, and KLD loss functions under different temperature parameters, with the experimental results shown in Table VII. The performance of these losses is compared based on the average results across the three splits in the NTU-60 dataset. The KLD loss demonstrates the best performance, confirming the effectiveness of using more positive samples for contrastive learning. In contrast, the InfoNCE and SoftmaxCE losses define the positive samples as only the corresponding semantic feature for each visual feature, resulting in fewer positive samples and thus suboptimal performance.
@@
V Conclusion
This work introduces a novel Dual Visual-Text Alignment (DVTA) for skeleton-based zero-shot action recognition. DVTA mainly comprises two alignment modules, Direct Alignment (DA) module and Augmented Alignment (AA) module, and the Semantic Description Enhancement (SDE) submodule. Our approach considerably outperforms existing zero-shot state-of-the-art methods on NTU RGB+D 60, NTU RGB+D 120 and PKU-MMD. Detailed ablation studies demonstrate the effectiveness of the three components as well as the validity of designed projectors and the LeakySigmoid activation function. Visualizations of similarity scores between skeleton features and text features confirm the complementarity between the DA and the Augmented Alignment. We believe that the proposed method can be applied to the general task of zero-shot or open-set action recognition. In the future, we will combine prompts to learn better alignment between visual and text modalities for zero-shot human action understanding.

### v2

@@ 3.2.2 Augmented Alignment
4 Experiments
4.1 Datasets
4.2 Implementation Details
For data preprocessing, we resize the skeleton sequence to 50 frames using linear interpolation and removal of invalid frames, which is the same as on Cross-CLR [40]. For text modality, 768-dimensional word embeddings are obtained by a pre-trained text feature extractor from Sentence-Bert. For skeleton modality, Shift-GCN [19] are used as feature encoders to obtain 256-dimensional features. The dimensionality is then increased to 768 dimensions by the nonlinear feature projector. The semantic features are passed through a linear layer to enhance the learning of similarity and to retain as much semantic information as possible. The optimizer employed is Adam, with a learning rate of 1e-5, and the training process spans 100 epochs, utilizing the CosineAnnealing scheduler. The mini-batch size is 128, and the temperature parameter for the contrastive loss is set to 0.1. Our implementation relies on the PyTorch framework, and all experiments are carried out using an NVIDIA GeForce RTX 4090.
4.3 Experimental Results
Comparison with State-of-the-arts: We compare the proposed DVTA with the state-of-the-art approaches to demonstrate its effectiveness in zero-shot learning. To ensure a fair comparison, our experimental setup strictly remains consistent with that of SynSE [27], including the use of identical skeleton feature and semantic feature encoders as well as the same split between seen and unseen classes. This setting incorporates only two datasets: NTU-60 and NTU-120. We employ Shift-GCN [19] as the skeleton feature extractor and leverage the pre-trained Sentence-Bert for semantic feature extraction. Follwing SynSE [27], we establish two distinct category divisions for each dataset. The NTU-60 dataset divided its 60 categories into 5 and 12 unseen classes, corresponding to the 55/5 and 48/12 splits in the table. For the NTU-120 dataset, which boasts a larger sample size and more categories, the number of unseen classes increased to 10 and 24, resulting in the 110/10 and 96/24 splits. Table 1 presents a comparison of the performance of the DVTA with state-of-the-art methods on the NTU-60 and NTU-120 datasets. The proposed DVTA achieves excellent performance in zero-shot learning. Compared with SMIE, DVTA achieves 1.3% and 3.96% performance improvement on the 55/5 split and 48/12 split of the NTU-60 dataset, respectively. For the NTU-120 dataset, DVTA achieves 9.15% and 6.51% on the 110/10 split and 96/24 split, respectively. “DVTA w/o SDE” means zero-shot learning on the pure network architecture without Semantic Description Enhancement. The performance of all settings exceeds the SMIE, reflecting the superiority of our network architecture. It is noteworthy that as the number of unseen classes increases, zero-shot learning becomes more challenging. However, our DVTA still performs remarkably well, especially on the more challenging 110/10 split and 96/24 split, which fully demonstrates the significant advantage of our model in handling scenarios with more unseen classes.
Table 1: Comparison of state-of-the-art methods with DVTA on the NTU-60 and NTU-120 datasets.
@@ 4.3 Experimental Results
Table 2: Optimized experimental setting on NTU-60, NTU-120, and PKU-MMD datasets.
@@
Optimized Experimental Setting Results: To better validate the effectiveness of our method in zero-shot learning, we conduct comparison experiments under an optimized setting [6]. Specifically, we use three large-scale skeleton datasets, i.e., NTU-60, NTU-120, and PKU-MMD as benchmark datasets. For each dataset, we conduct three randomized splits and average the results to reduce biases introduced by different class splits. The seen/unseen class splits are set as 55/5, 110/10 and 46/5, respectively, following the class split settings in SMIE [6]. To better focus on evaluating the effectiveness of the alignment model, we adopt the classical ST-GCN [13] as the visual feature extractor to minimize the impact of complex feature extractor architectures on the results. Sentence-BERT is used for semantic feature extraction and evaluation. The comparison results are shown in Table 2. Our method outperforms others across all three datasets, achieving improvements of 10.46%, 3.96%, and 9.91%, respectively. The results indicate that DVTA effectively aligns the skeleton and text modalities, and the dual-alignment approach is more conducive to zero-shot learning.
Table 3: Influence of different modules on the NTU-60 and NTU-120 datasets.
@@
4.4 Ablation Study and Visualizations
Influence of Modules: To evaluate the effectiveness of each module in the Dual Visual-Text Alignment (DVTA), we conduct a systematic ablation study. The experiments are performed under the optimized settings, with three split tests on the NTU-60 and NTU-120 datasets, and the average performance is reported, as shown in Table 3.
  • 1.
  • 2.
  • 3.
  • 4.
Table 4: Effect on different settings of Direct Alignment module.
@@ 4.4 Ablation Study and Visualizations
The DA module can be applied as an independent model for zero-shot skeleton action recognition, optimized through training with the $L_{1}$ loss function. This module comprises a deep nonlinear mapping projector for skeleton features and a linear layer for text features, allowing for the investigation of the utility of these components through ablation experiments. The results of the ablated experiments on the DA module’s components are shown in Table 4, where “Semantic” denotes the use of the text mapping, and “Visual” denotes the use of the designed deep nonlinear mapping projector for skeleton features. The results indicate that using the complete DA module more effectively reduces the gap between modalities, thereby improving alignment quality.
Table 5: Effect on different settings of the hyperparameter $\gamma$ on Augmented Alignment module.
@@
Effect of Augmented Alignment module: In the Augmented Alignment (AA) module, we design a LeakySigmoid function to convert the output value of the network $E$ into a similarity score $G$. It has a hyperparameter, $\gamma$, which controls the slope of the function at negative values. To determine the optimal choice of $\gamma$ and demonstrate the effectiveness of the LeakySigmoid function, we conducted experiments on all four splits based on the complete model. The experimental results are shown in Table 5, where “None” indicates no utilization of any transformation function, and “Sigmoid” signifies the application of the Sigmoid function to output similarity scores. Through comparison, it is found that the performance of the LeakySigmoid function consistently surpasses that of the Sigmoid function, regardless of the value set for the $\gamma$ hyperparameter. This discovery effectively validates the usefulness of the LeakySigmoid function in the AA module. Specifically, when $\gamma$ was set to 0.01, the model exhibits the best performance on two splits. To ensure consistency and comparability in our experiments, we uniformly set the $\gamma$ hyperparameter to 0.01 throughout the entire experimental process.
Table 6: Ablation on Semantic Descriptions Enhancement module.
@@
Effect of Semantic Descriptions Enhancement module: We encode the action text using different text encoders. Table 6 presents the SDE ablation study results using Sentence-BERT and CLIP as semantic feature extractors on the NTU-60 and NTU-120 datasets. The results indicate that the use of the Semantics Description Enhancement (SDE) module leads to significant performance improvements across all split tests, regardless of the encoder used. For the NTU-120 dataset, under the 96/24 split, the accuracy of action labels and the accuracy when enhanced with action description encoding are 48.53% and 51.81%, respectively, using the BERT encoder. Figure 6 shows the per-class accuracy changes after applying the SDE module. The visualization reveals substantial improvements in the accuracy of certain hand-related action categories, such as “waving”, “picking up”, and “punching and kicking”.
Table 7: Performance of different skeleton and text encoders on the NTU-60 and NTU-120 datasets.
Skeleton Encoder Text Encoder NTU-60(%) NTU-120(%)
55/5 48/12 110/10 96/24
ST-GCN [13] BERT 73.99 41.32 63.22 47.99
ST-GCN [13] CLIP 75.51 43.37 61.91 44.54
Shift-GCN [19] BERT 79.28 44.14 74.89 51.81
Shift-GCN [19] CLIP 79.64 47.48 64.11 48.18
CTR-GCN [21] BERT 81.07 45.92 76.86 54.70
CTR-GCN [21] CLIP 82.55 48.40 68.26 56.97
Effect of Different Encoders: Table 7 reports the performance of our proposed method on the NTU-60 and NTU-120 datasets using different visual and text encoders. We adopt several widely used skeleton-based backbones as visual encoders, including ST-GCN, Shift-GCN and CTR-GCN. Sentence-BERT and CLIP are used as text encoders for semantic feature extraction. The results show that performance varies across different text encoders, possibly due to the diverse knowledge embedded in them, some of which may not directly relate to action class semantics. Nevertheless, irrespective of the text encoder used, our method consistently achieves better performance when equipped with more advanced visual feature extractors. This suggests that the proposed approach exhibits stronger zero-shot learning capabilities when combined with more sophisticated skeleton backbones.
Table 8: Experiments with the loss function on the NTU-60 dataset with different temperature parameters.
@@
Effect of the Loss Functions: We evaluate the InfoNCE, SoftmaxCE, and KLD loss functions under different temperature parameters, with the experimental results shown in Table 8. The performance of these losses is compared based on the average results across the three splits in the NTU-60 dataset. The KLD loss demonstrates the best performance, confirming the effectiveness of using more positive samples for contrastive learning. In contrast, the InfoNCE and SoftmaxCE losses define the positive samples as only the corresponding semantic feature for each visual feature, resulting in fewer positive samples and thus suboptimal performance.
Table 9: Computational efficiency comparison. Params, FLOPs, and FPS represent the number of learnable parameters, the floating point operations per sample, and the number of samples processed per second during inference, respectively.
Method Params FLOPs FPS NTU-60
SMIE [6] 1.57M 3.14M 1623 40.18
PURLS [32] 7.58M 17.32M 497 40.99
DVTA 4.88M 4.87M 1335 44.14
Computational Efficiency Analysis: As shown in Table 9, we compare the Params, FLOPs, and FPS of our method (DVTA) with SMIE and PURLS, evaluated on the same hardware setup (NVIDIA GeForce RTX 4090). For these metrics, we measure the computational efficiency of the alignment model independently, using pre-extracted visual and text features as input. Inference time is measured on the NTU-60 dataset of 110/10 split by running the models on the same number of test samples, and FPS is computed by dividing the total inference time by the number of samples. The results indicate that both SMIE and DVTA maintain high computational efficiency. PURLS exhibits lower efficiency due to its extensive use of attention mechanisms. Overall, DVTA achieves competitive performance while maintaining relatively low computational complexity.
@@
Figure 8: The Zero-shot Recognition Results on the 55/5 split of the NTU-60 dataset.
Qualitative Results and Analysis: Figure 8 illustrates the skeleton visualization of the 55/5 split on the NTU-60 dataset. The left side showcases the skeleton dynamics of five unseen class actions, while the right side displays the action prediction scores. The action ”put on a hat”, which involves raising the hands above the head, is similar to ”jump up”, however, our model makes the correct prediction. Furthermore, ”jump up” ranking as the second highest prediction score also indicates that our model has reasonable predictions for similar actions. The prediction scores for other actions also support the effectiveness of the method.
5 Conclusion
This work introduces a novel Dual Visual-Text Alignment (DVTA) for skeleton-based zero-shot action recognition. DVTA mainly comprises two alignment modules, Direct Alignment (DA) module and Augmented Alignment (AA) module, and the Semantic Description Enhancement (SDE) submodule. Our approach considerably outperforms existing zero-shot state-of-the-art methods on NTU RGB+D 60, NTU RGB+D 120 and PKU-MMD. Detailed ablation studies demonstrate the effectiveness of the three components as well as the validity of designed projectors and the LeakySigmoid activation function. Visualizations of similarity scores between skeleton features and text features confirm the complementarity between the DA and the Augmented Alignment. We believe that the proposed method can be applied to the general task of zero-shot or open-set action recognition. However, a limitation of our approach is its reliance on pre-extracted features. In the future, we will focus on learning joint visual-text representations end-to-end, allowing for more dynamic feature extraction and improving the applicability of the method to real-time and large-scale scenarios.
Acknowledgement
This research was supported by the Jiangsu Science Foundation (BK20230833), the National Science Foundation of China (62302093, 62172090), Start-up Research Fund of Southeast University under Grant RF1028623097, and the Big Data Computing Center of Southeast University.